# Searching for blue in the dark

### v1

@@ IV Methodology
In this section we describe our procedure for forecasting the sensitivity of future 21-cm LIM experiments to enhancements of small-scale power. We consider three proposed future instruments as examples: one Earth-based instrument, inspired by the Square Kilometre Array Observatory (SKAO), and two lunar instruments with different sensitivities.
For our forecasts, we take $\Lambda$CDM as a reference model, using the Planck 2018 best fit values Aghanim et al. (2020) for the fiducial cosmological parameters: $\{\omega_{b}=0.02238,\ \omega_{c}=0.1201,\ h=0.6781,\ n_{s}=0.96605,\ \ln 10^{10}A_{s}=3.0448\}$.
@@ IV.1 Forecasting setup
We estimate the detectability of a given set of parameters controlling the deviation from $\Lambda$CDM of the primordial power spectrum by calculating the signal-to-noise ratio (SNR) as Scelfo et al. (2018)
@@
We model the noise power spectrum as Shiraishi et al. (2016)
@@
where $t_{\rm obs}$ is the total time of observation, and $f_{\rm cover}$ is the coverage fraction (i.e., the fraction of the array area that is covered by antennas). We take the system temperature to be the synchrotron temperature of the observed sky Shiraishi et al. (2016):
@@
We assume that the redshift bins are independent, which is a good approximation provided that the correlation length in frequency is smaller than the bandwidth of the instrument. We define the radial correlation length $\xi_{r}$ as the radial separation, beyond which the cross-correlation between two redshift slices is less than 1/2 of the power spectrum Muñoz et al. (2015). The corresponding correlation length in frequency is
@@
We take linearly spaced bins in frequency, which we convert into redshift bins, and verify that $\xi_{\nu}$ is smaller than the bandwith $B$ of the instrument. We neglect cross-correlations between different redshift bins, since they are expected to be negligible for 21-cm LIM measurements Hall et al. (2013).
For the case of running of the spectral tilt, we perform a Fisher matrix analysis FISHER (1935); Coe (2009); Bellomo et al. (2020); Bernal et al. (2020); Verde (2010); Heavens (2010); Tegmark et al. (1997); Vogeley and Szalay (1996); Tegmark et al. (2004); Seo and Eisenstein (2003), since we have a specific model with fiducial values for the running parameters ($\alpha_{s}$, $\beta_{s}$); whereas for the blue-tilted curvature and isocurvature scenarios, we use the SNR to investigate which values of $\{k_{b},n_{s}\}$ or $\{A_{\rm iso},n_{\rm iso}\}$ we can test. The Fisher approach approximates the likelihood around its maximum as a Gaussian and returns the smallest possible error on the parameter, namely the one set by the Cramer-Rao bound. The Fisher information matrix is
@@ IV.2 Instruments
There are several proposals for low-frequency interferometers that would be able to detect the redshifted 21-cm signal from the dark ages. Most proposals are for lunar-based experiments, see e.g., Burns et al. (2017); Burns et al. (2019a); Burns et al. (2019b); Bale et al. (2023); Chen et al. (2023); Polidan et al. (2024); Lazio et al. (2009); Price et al. (2018); Goel et al. (2022), including lunar-orbiting CubeSats Artuc and Acedo (2024); more proposals are being developed for both lunar- and space-based interferometers Raccanelli et al. (in prep.). These experiments are generally lunar-based, because the frequency of the redshifted signal from the end of the dark ages lies at the edge of where the Earth’s ionosphere becomes opaque Jester and Falcke (2009); Datta et al. (2016). Moreover, the far side of the Moon should be completely radio quiet and stable Koopmans et al. (2021), making it an ideal place for such an instrument.
There is an open debate on how well we will be able to model the signal and actually observe the end of the dark ages from the (Earth) ground. Several Earth-based instruments had planned to observe the dark ages Bevins et al. (2022); Jishnu Nambissan et al. (2021); Blyth et al. (2015); Maio et al. (2015); Monsalve et al. (2024), but this possibility is still uncertain; therefore, in this work we also consider an Earth-based instrument, assuming it will observe the very end of the dark ages, which we take to be at redshift $z=30$.
In order to keep our findings as general as possible, we consider three surveys, the details of which are presented in Table 1. The instruments are a future Earth-based instrument—an advanced version of the SKAO, which we refer to as SKAO-like, limited to $z=30$—and two possible configurations of a lunar radio array (LRA) on the far side of the Moon. We refer to the lunar instruments as LRAI and LRAII; we consider a baseline of 100 km for the former, while the latter has a baseline that spans nearly the diameter of the Moon to demonstrate the maximal capabilities of a lunar array.
@@ V Results
Here we present our forecasted constraints on the models presented in Section II, using the methodology of Section IV.1 for the instruments of Section IV.2.
@@ V.1 Enhanced curvature power spectrum
We start by showing results for the parameterization in Eq. (2). In Fig. 4 we show the SNR for the detection of deviations from $\Lambda$CDM, as a function of the cutoff scale and slope $\{k_{b},n_{b}\}$, for each instrument considered in this work. In the $k$ range that is accessible to a given instrument, virtually any blue tilts would be easily detected with a high SNR. We expect this result, given the error bars shown in Fig. 3. Importantly, the baseline of the interferometer is a key feature for testing small-scale enhancements of the power spectrum. This point is also illustrated in Fig. 5, where we show how the SNR changes as a function of $k_{b}$ for specific values of $n_{b}=2,3,4$.
@@
The highest value of $k_{b}$ that can be probed is nearly identical to the highest $k$ measured by the instrument, $k_{\rm max}=\ell_{\rm max}\times\chi(z)$, where $\ell_{\rm max}$ is given in Eq. (20) and $\chi(z)$ is the comoving distance. At $z=30$, $k_{\rm max}\sim 8\,{\rm Mpc}^{-1}$ for the small instruments (SKAO-like, LRAI), while $k_{\rm max}\sim 230\,{\rm Mpc}^{-1}$ for the larger instrument (LRAII). Only for the lowest values of $n_{b}$ and the highest $k_{b}$ is the signal indistinguishable from the noise.
@@ V.2 Enhanced dark matter isocurvature power spectrum
We now consider cold dark matter isocurvature modes that produce a blue-tilted isocurvature primordial power spectrum, given by Eq. (3). Figure 6 shows the SNR for the detection of isocurvature modes, as a function of the amplitude ratio $A_{\rm iso}/A_{s}$ and the spectral tilt $n_{\rm iso}$, for each instrument considered in this work. The largest value we show for $A_{\rm iso}/A_{s}$ coincides with the upper bound from Planck Akrami et al. (2020). An Earth-based instrument tapping into the dark ages can improve on the Planck constraints by about an order of magnitude, depending on the spectral index. The smaller lunar array, LRAI, would significantly increase the significance of such constraints, while the larger LRAII could improve the constraints by two orders of magnitude for very low value of $n_{\rm iso}$ and several orders more for higher tilts.
The sensitivity of 21-cm dark ages measurements would be better than that for future galaxy surveys. A large scale structure survey like Euclid allows, at best, to detect ($A_{\rm iso}/A_{s}=0.09,n_{\rm iso}=3$), while using an idealized version of MegaMapper could detect ($A_{\rm iso}/A_{s}=0.015,n_{\rm iso}=3$; $A_{\rm iso}/A_{s}=0.002,n_{\rm iso}=4$Chung et al. (2023).
@@ V.3 Running parameters
We perform a Fisher forecast on the standard cosmological parameters $\{\omega_{b},\omega_{c},h,n_{s},\ln 10^{10}A_{s}\}$, plus the running parameters $\{\alpha_{s},\beta_{s}\}$. For fiducial values we take those of single-field slow-roll (SFSR) inflation: $\{\alpha_{s},\beta_{s}\}=\{10^{-3},10^{-5}\}$ Kosowsky and Turner (1995); Cabass et al. (2016). We then perform two analyses: one varying all the parameters above and one fixing cosmological ones to their Planck 2018 best fit values Aghanim et al. (2020).
@@
As a point of comparison, Ref. Muñoz et al. (2017) has forecasts for the spectral running parameters for future CMB and LSS experiments. While an Earth-based interferometer would be competitive with future CMB measurements, it would not provide strong improvements. The lunar arrays, however, would be more sensitive than future CMB+galaxy surveys, especially for the $\beta_{s}$ parameter, for which the improvement could be orders of magnitude.
@@ VI Conclusions
The primordial power spectrum is well-constrained by CMB and LSS data to be almost scale-invariant (and slightly red-tilted) at large scales. Several new-physics and dark matter models can imprint small-scale deviations from $\Lambda$CDM and are still largely unconstrained. Many of these models induce an increase of power at smaller scales, producing a blue-tilted power spectrum.
@@ Acknowledgements.
AR thanks the University of Texas, Austin, for the hospitality during the initial stages of this work. KB thanks the University of Padova for the hospitality during the intermediate stages of this work. We thank Mustafa Amin, Nicola Bellomo, Andrew Long, and Julian Muñoz for useful discussions. KB acknowledges support from the National Science Foundation (NSF) under Grant No. PHY-2112884 and acknowledges the “Dark Matter Theory, Simulation, and Analysis in the Era of Large Surveys” workshop and the Kavli Institute for Theoretical Physics (KITP) for the hospitality and support under NSF Grant No. PHY-2309135. AR acknowledges funding from the Italian Ministry of University and Research (MIUR) through the “Dipartimenti di eccellenza” project “Science of the Universe”.

### v2

@@ IV Methodology
In this section we describe our procedure for forecasting the sensitivity of future 21-cm LIM experiments to enhancements of small-scale power. We consider three proposed future instruments as examples: one Earth-based instrument, inspired by the Square Kilometre Array Observatory, and two lunar instruments with different sensitivities.
For our forecasts, we take $\Lambda$CDM as a reference model, using the Planck 2018 best fit values [1] for the fiducial cosmological parameters: $\{\omega_{b}=0.02238,\ \omega_{c}=0.1201,\ h=0.6781,\ n_{s}=0.96605,\ \ln 10^{10}A_{s}=3.0448\}$.
@@ IV.1 Forecasting setup
We estimate the detectability of a given set of parameters controlling the deviation from $\Lambda$CDM of the primordial power spectrum by calculating the signal-to-noise ratio (SNR) as [127]
@@
We model the noise power spectrum as [128]
@@
where $t_{\rm obs}$ is the total time of observation, and $f_{\rm cover}$ is the coverage fraction (i.e., the fraction of the array area that is covered by antennas). We emphasize Eq. (21) is the noise power spectrum of the instrument. In this work we do not consider possible additional sources of noise such as foregrounds, lensing, or atmospheric noise for the Earth-based instrument. We take the system temperature to be the synchrotron temperature of the observed sky [128]:
@@
We assume that the redshift bins are independent, which is a good approximation provided that the correlation length in frequency is smaller than the bandwidth of the instrument. We define the radial correlation length $\xi_{r}$ as the radial separation, beyond which the cross-correlation between two redshift slices is less than 1/2 of the power spectrum [129]. The corresponding correlation length in frequency is
@@
We take linearly spaced bins in frequency, which we convert into redshift bins, and verify that $\xi_{\nu}$ is smaller than the bandwith $B$ of the instrument. We neglect cross-correlations between different redshift bins, since they are expected to be negligible for 21-cm LIM measurements [130].
For the case of running of the spectral tilt, we perform a Fisher matrix analysis [131, 132, 133, 134, 135, 136, 137, 138, 139, 140], since we have a specific model with fiducial values for the running parameters ($\alpha_{s}$, $\beta_{s}$); whereas for the blue-tilted curvature and isocurvature scenarios, we use the SNR to investigate which values of $\{k_{b},n_{s}\}$ or $\{A_{\rm iso},n_{\rm iso}\}$ we can test. The Fisher approach approximates the likelihood around its maximum as a Gaussian and returns the smallest possible error on the parameter, namely the one set by the Cramer-Rao bound. The Fisher information matrix is
@@ IV.2 Instruments
There are several proposals for low-frequency interferometers that would be able to detect the redshifted 21-cm signal from the dark ages. Most proposals are for lunar-based experiments, see e.g., [70, 71, 72, 73, 74, 75, 76, 77, 141], including lunar-orbiting CubeSats [78]; more proposals are being developed for both lunar- and space-based interferometers [142]. These experiments are generally lunar-based, because the frequency of the redshifted signal from the end of the dark ages lies at the edge of where the Earth’s ionosphere becomes opaque [143, 144]. Moreover, the far side of the Moon should be completely radio quiet and stable [145], making it an ideal place for such an instrument.
There is an open debate on how well we will be able to model the signal and actually observe the end of the dark ages from the (Earth) ground. Several Earth-based instruments had planned to observe the dark ages [79, 80, 81, 82, 83], but this possibility is still uncertain; therefore, in this work we also consider an Earth-based instrument, assuming it will observe the very end of the dark ages, which we take to be at redshift $z=30$. We consider an idealized scenario where the signal is only affected by the noise of the instrument itself; the modeling of the noise from the Earth’s atmosphere is beyond the scope of this work.
In order to keep our findings as general as possible, we consider three surveys, the details of which are presented in Table 1. The instruments are an idealized future Earth-based instrument—chosen to be of similar size as SKAO, which we therefore refer to as SKAO-like, limited to $z=30$—and two possible configurations of a lunar radio array (LRA) on the far side of the Moon. We refer to the lunar instruments as LRAI and LRAII; we consider a baseline of 100 km for the former, while the latter has a baseline that spans nearly the diameter of the Moon to demonstrate the maximal capabilities of a lunar array. We also consider how our results are affected by changes to the total observation time $t_{\rm obs}$ and the coverage fraction $f_{\rm cover}$ in Appendix A. The specific distribution of the radio arrays can also affect the sensitivity of the instrument [146, 147].
@@ V Results
Here we present our forecasted constraints on the models presented in Sec. II, using the methodology of Sec. IV.1 for the instruments of Sec. IV.2.
@@ V.1 Enhanced curvature power spectrum
We start by showing results for the parametrization in Eq. (2). In Fig. 4 we show the SNR for the detection of deviations from $\Lambda$CDM, as a function of the cutoff scale and slope $\{k_{b},n_{b}\}$, for each instrument considered in this work. In the $k$ range that is accessible to a given instrument, virtually any blue tilts would be easily detected with a high SNR. We expect this result, given the error bars shown in Fig. 3. Importantly, the baseline of the interferometer is a key feature for testing small-scale enhancements of the power spectrum. This point is also illustrated in Fig. 5, where we show how the SNR changes as a function of $k_{b}$ for specific values of $n_{b}=2,3,4$.
@@
The highest value of $k_{b}$ that can be probed is nearly identical to the highest $k$ measured by the instrument, $k_{\rm max}=\ell_{\rm max}/\chi(z)$, where $\ell_{\rm max}$ is given in Eq. (20) and $\chi(z)$ is the comoving distance. At $z=30$, $k_{\rm max}\sim 8\,{\rm Mpc}^{-1}$ for the small instruments (SKAO-like, LRAI), while $k_{\rm max}\sim 230\,{\rm Mpc}^{-1}$ for the larger instrument (LRAII). Only for the lowest values of $n_{b}$ and the highest $k_{b}$ is the signal indistinguishable from the noise.
@@ V.2 Enhanced dark matter isocurvature power spectrum
We now consider cold dark matter isocurvature modes that produce a blue-tilted isocurvature primordial power spectrum, given by Eq. (3). Figure 6 shows the SNR for the detection of isocurvature modes, as a function of the amplitude ratio $A_{\rm iso}/A_{s}$ and the spectral tilt $n_{\rm iso}$, for each instrument considered in this work. The largest value we show for $A_{\rm iso}/A_{s}$ coincides with the upper bound from Planck [7]. An Earth-based instrument tapping into the dark ages can improve on the Planck constraints by about an order of magnitude, depending on the spectral index. The smaller lunar array, LRAI, would significantly increase the significance of such constraints, while the larger LRAII could improve the constraints by two orders of magnitude for very low value of $n_{\rm iso}$ and several orders more for higher tilts.
The sensitivity of 21-cm dark ages measurements would be better than that for future galaxy surveys. A large scale structure survey like Euclid allows, at best, to detect ($A_{\rm iso}/A_{s}=0.09,n_{\rm iso}=3$), while using an idealized version of MegaMapper could detect ($A_{\rm iso}/A_{s}=0.015,n_{\rm iso}=3$; $A_{\rm iso}/A_{s}=0.002,n_{\rm iso}=4$[61].
@@ V.3 Running parameters
We perform a Fisher forecast on the standard cosmological parameters $\{\omega_{b},\omega_{c},h,n_{s},\ln 10^{10}A_{s}\}$, plus the running parameters $\{\alpha_{s},\beta_{s}\}$. For fiducial values we take those of single-field slow-roll (SFSR) inflation: $\{\alpha_{s},\beta_{s}\}=\{10^{-3},10^{-5}\}$ [104, 106]. We then perform two analyses: one varying all the parameters above and one fixing cosmological ones to their Planck 2018 best fit values [1].
@@
As a point of comparison, Ref. [89] has forecasts for the spectral running parameters for future CMB and LSS experiments. While an Earth-based interferometer would be competitive with future CMB measurements, it would not provide strong improvements. The lunar arrays, however, would be more sensitive than future CMB+galaxy surveys, especially for the $\beta_{s}$ parameter, for which the improvement could be orders of magnitude.
@@ VI Conclusions
The primordial power spectrum is well-constrained by CMB and LSS data to be almost scale invariant (and slightly red tilted) at large scales. Several new-physics and dark matter models can imprint small-scale deviations from $\Lambda$CDM and are still largely unconstrained. Many of these models induce an increase of power at smaller scales, producing a blue-tilted power spectrum.
@@ Acknowledgements.
A.R. thanks the University of Texas, Austin, for the hospitality during the initial stages of this work. K.B. thanks the University of Padova for the hospitality during the intermediate stages of this work. We thank Mustafa Amin, Sabino Matarrese, Nicola Bellomo, Andrew Long, and Julian Muñoz for useful discussions. K.B. acknowledges support from the National Science Foundation (NSF) under Grant No. PHY-2112884 and acknowledges the “Dark Matter Theory, Simulation, and Analysis in the Era of Large Surveys” workshop and the Kavli Institute for Theoretical Physics (KITP) for the hospitality and support under NSF Grant No. PHY-2309135. A.R. acknowledges funding from the Italian Ministry of University and Research (MIUR) through the “Dipartimenti di eccellenza” project “Science of the Universe”.
Appendix A SNR dependence on instrument specifications
In this Appendix, we consider how the total observational time $t_{\rm obs}$ and the fraction of the array area covered by antennas $f_{\rm cover}$ affects the SNR for SKAO-like (Fig. 7) and LRAII (Fig. 8). For both instruments, it is clear that even with a much shorter running time, the obtained signal is still very strong. Figure 7 shows that for SKAO-like, we need at least $f_{\rm cover}=0.2$ in order not to lose sensitivity on the smallest scales. Similarly for LRAII in Fig. 7, we require $f_{\rm cover}\gtrsim 0.3$ to probe the smallest scales.
Figure 7: SNR as a function of $k_{b}$ for a blue-tilted deviation from $\Lambda$CDM, using the SKAO-like instrument as specified in Table 1 but with different values for $f_{\rm cover}$ (top) and $t_{\rm obs}$ (bottom), as indicated in the captions. We set $n_{b}=3$ and indicate $\rm{SNR}=5$ with a gray horizontal line for reference.
Figure 8: SNR as a function of $k_{b}$ for a blue-tilted deviation from $\Lambda$CDM, using the LRAII instrument as specified in Table 1 but with different values for $f_{\rm cover}$ (top) and $t_{\rm obs}$ (bottom), as indicated in the captions. We set $n_{b}=3$ and indicate $\rm{SNR}=5$ with a gray horizontal line for reference.